# All meromorphic solutions of Fermat-type functional equations

### v1

@@ 1 Introduction and main results
which implies that the above equation admits meromorphic solution, which is not the form $de^{\frac{\alpha z+\beta}{3}}$. Obviously, the hyper-order of $f$ is $\rho_{2}(f)=1$.
@@
Obviously, all the points in $\mathbb{L}$ are the poles and periods of Weierstrass-$\wp$. Suppose that $D$ is the parallelogram with vertices at 0, $\omega_{1}$, $\omega_{2}$, $\omega_{1}+\omega_{2}$. Note that the order of Weierstrass-$\wp$ is 2. Here, the order is the number of poles of Weierstrass-$\wp$ or the number of zeros of $\wp-a$ ($a\in\mathbb{C}$) in the parallelogram. Together with $(\wp^{\prime})2=4\wp^{3}-1$, one gets that Weierstrass-$\wp$ has two distinct zeros in $D$, say $\theta_{1}$ and $\theta_{2}$. In view of $\wp(z)=0\Leftrightarrow\wp^{\prime}(z)=\pm i$, without loss of generality, throughout the paper, we assume that $\wp^{\prime}(\theta_{1})=-i$ and $\wp^{\prime}(\theta_{2})=i$.
@@
(1). $h(L(z))=-h(z)+\theta_{1}+\tau_{1}$ with $\tau_{1}\in\mathbb{L}$;
(2). $h(L(z))=-h(z)+\theta_{2}+\tau_{2}$ with $\tau_{2}\in\mathbb{L}$;
(3). $h(L(z))=-h(z)+\tau_{3}$ with $\tau_{3}\in\mathbb{L}$.
@@
(1). $h(z+c)=-h(z)+\theta_{1}+\tau_{1}$ with $\tau_{1}\in\mathbb{L}$;
(2). $h(z+c)=-h(z)+\theta_{2}+\tau_{2}$ with $\tau_{2}\in\mathbb{L}$;
(3). $h(z+c)=-h(z)+\tau_{3}$ with $\tau_{3}\in\mathbb{L}$.
@@
Example 2. In [39, Theorem 3], Ozawa derived that there exists periodic entire function of arbitrary order $\sigma(\geq 1)$. So, there exists an entire function $g(z)$ such that $\rho(g(z))=\sigma$ and $g(z+c)=g(z)$. Set $h(z)=e^{az}g(z)+b$, where $a$, $b$ are constants with $e^{ac}=-1$ and $2b=\theta_{1}+\tau_{1}$ or $\theta_{2}+\tau_{2}$ or $\tau_{3}$. Then, $\rho(h(z))=\rho(g(z))=\sigma$ and a calculation yields that
$h(z+c)=-h(z)+2b.$
@@
if and only if $f(z)=Ae^{\frac{g(z)}{n}}$ and $g(z+c)=\frac{n}{m}g(z)+nLn\frac{B}{A}$, where $A,~{}B$ are two nonzero constants with $A^{n}+B^{m}=1$.
In particularly, if the order of $g(z)$ is less than 1, then (1.9) admits nonconstant meromorphic solutions if and only if $n=m$ and $g(z)=\alpha z+\beta$, $f(z)=Ae^{\frac{\alpha z+\beta}{n}}$ with constant $\alpha(\neq 0)$ such that $A^{n}(1+e^{\alpha c})=1$.
@@
Remark 3. From Theorems 3, we see that the equation (1.9) do not admit meromorphic solutions with poles. It is pointed out that the same argument in Theorems 3 can deal with the Fermat-type difference equations $f^{m}(z)+f^{n}(z+c)=e^{g(z)}$ with $n(\geq 2),~{}m(\geq 3)$ and $(m,n)\neq(3,3)$, we omit the details here.
Combining all the theorems, we can derive Theorem 4.
Theorem 4. Suppose that $P(z)$ is a polynomial and $f(z)$ is a nonconstant meromorphic function. Suppose that $n(\geq 2),~{}m(\geq 2)$ are two integers with $m+n\geq 5$. Then, $f(z)$ is a solution of
@@
The following corollary immediately follows from Theorem 4.
Corollary. Suppose that $f(z)$ is a nonconstant meromorphic function, and $n(\geq 2),~{}m(\geq 2)$ are two integers with $m+n\geq 5$. Then, $f(z)$ is a solution of
@@ 2 Proofs of main results
Firstly, we will prove the necessity. Assume that $h(z)$ satisfies one of (1)-(3), we will prove that $f(z)$ is a meromorphic solution of (1.4). Suppose that $h(z)$ satisfies (1). Then $h(L(z))=-h(z)+\theta_{1}+\tau_{1}$ with $\tau_{1}\in\mathbb{L}$. It is know that
@@
$\wp(w-\theta_{1})$ $=\frac{1}{4}[\frac{\wp^{\prime}(w)-\wp^{\prime}(-\theta_{1})}{\wp(w)-\wp(-\theta_{1})}]^{2}-\wp(w)-\wp(-\theta_{1})=\frac{1}{4}[\frac{\wp^{\prime}(w)-i}{\wp(w)}]^{2}-\wp(w)$ (2.1)
$=\frac{1}{4}\frac{(\wp^{\prime}(w)-i)^{2}\wp(w)}{\wp(w)^{3}}-\wp(w)=\frac{(\wp^{\prime}(w)-i)^{2}\wp(w)}{\wp^{\prime}(w)^{2}+1}-\wp(w)$
$=\frac{(\wp^{\prime}(w)-i)^{2}\wp(w)}{(\wp^{\prime}(w)+i)(\wp^{\prime}(w)-i)}-\wp(w)=\frac{-2i\wp(w)}{\wp^{\prime}(w)+i}.$
$\wp^{\prime}(w-\theta_{1})$ $=(\frac{-2i\wp(w)}{\wp^{\prime}(w)+i})^{\prime}=-2i\frac{\wp^{\prime}(w)(\wp^{\prime}(w)+i)-\wp(w)\wp^{\prime\prime}(w)}{(\wp^{\prime}(w)+i)^{2}}$ (2.2)
$=-2i\frac{\wp^{\prime}(w)^{2}+i\wp^{\prime}(w)-\wp(w)\wp^{\prime\prime}(w)}{(\wp^{\prime}(w)+i)^{2}}=-2i\frac{\wp^{\prime}(w)^{2}+i\wp^{\prime}(w)-\frac{3}{2}(\wp^{\prime}(w)^{2}+1)}{(\wp^{\prime}(w)+i)^{2}}$
$=i\frac{\wp^{\prime}(w)-3i}{\wp^{\prime}(w)+i}.$
Set $h(z)=\omega$. Then $h(L(z))=-h(z)+\theta_{1}+\tau_{1}=-w+\theta_{1}+\tau_{1}$. Note that both Weierstrass-$\wp$ and $\wp^{\prime}$ are elliptic functions with periods $\omega_{1}$ and $\omega_{2}$. Together with the fact Weierstrass-$\wp$ is an even function and $\wp^{\prime}$ is an odd function, we have
$\wp(h(L(z)))=\wp(-w+\theta_{1}+\tau_{1})=\wp(w-\theta_{1}),$
$\wp^{\prime}(h(L(z)))=\wp^{\prime}(-w+\theta_{1}+\tau_{1})=-\wp^{\prime}(w-\theta_{1}).$
@@
$f(L(z))$ $=\frac{1}{2}\frac{1+\frac{\wp^{\prime}(h(L(z)))}{\sqrt{3}}}{\wp(h(L(z)))}=\frac{1}{2}\frac{1-\frac{\wp^{\prime}(w-\theta_{1})}{\sqrt{3}}}{\wp(w-\theta_{1})}=\frac{1}{2}\frac{1-\frac{1}{\sqrt{3}}i\frac{\wp^{\prime}(w)-3i}{\wp^{\prime}(w)+i}}{\frac{-2i\wp(w)}{\wp^{\prime}(w)+i}}$ (2.3)
$=\frac{1}{2}\frac{(\wp^{\prime}(w)+i)-\frac{1}{\sqrt{3}}i(\wp^{\prime}(w)-3i)}{-2i\wp(w)}=\frac{1}{2}\frac{-1-\sqrt{3}i}{2}\frac{1-\frac{\wp^{\prime}(w)}{\sqrt{3}}}{\wp(w)}=\frac{\eta}{2}\frac{1-\frac{\wp^{\prime}(w)}{\sqrt{3}}}{\wp(w)},$
where $\eta=\frac{-1-\sqrt{3}i}{2}$ and $\eta^{3}=1$. Thus, $f(L(z))$ has the form of $g(z)$ in (i) of Theorem A, which implies that $f(z)^{3}+f(L(z))^{3}=1$ and $f(z)$ is a solution of (1.4). If $h(z)$ satisfies (2) or (3), then the same argument yields that $f(z)$ is also a solution of (1.4).
@@
Observe that $A=\frac{(1+i\frac{\sqrt{3}}{3})}{\eta(1-i\frac{\sqrt{3}}{3})}$ and $\eta^{3}=1$. A calculation yields
@@
Suppose that $A=\frac{1}{2}+\frac{\sqrt{3}}{2}i=e^{\frac{\pi}{3}i}$. Among all the periods of $\wp(\omega)$, we list the periods which may take the smallest modulus as follows: $\omega_{1}$, $\omega_{1}+\omega_{2}$, $\omega_{2}$, $\omega_{2}-\omega_{1}$, $-\omega_{1}$, $-\omega_{2}-\omega_{1}$, $-\omega_{2}$ and $\omega_{1}-\omega_{2}$. Then, in view of that $A\tau=\tau e^{\frac{\pi}{3}i}$ is also a period of Weierstrass-$\wp$, we have two possible subcases (see Figures 1-2 below).
Figure 1: Subcase 2.1
Figure 2: Subcase 2.2
Subcase 2.1. $A\omega_{1}=\omega_{2}$ and $A\omega_{2}=\omega_{2}-\omega_{1}$ and $A(\omega_{2}-\omega_{1})=-\omega_{1}$.
Then, we have
$A^{2}\omega_{1}=A(A\omega_{1})=A\omega_{2}=A(\omega_{2}-\omega_{1})=-\omega_{1},$
which yields $A=\pm i$, a contradiction.
Subcase 2.2. $A\omega_{1}=\omega_{1}+\omega_{2}$ and $A(\omega_{1}+\omega_{2})=\omega_{2}$ and $A\omega_{2}=-\omega_{1}$.
We also have
$A^{2}\omega_{1}=A(\omega_{1}+\omega_{2})=A\omega_{2}=-\omega_{1},$
which also yields $A=\pm i$, a contradiction.
So, the case $A=\frac{1}{2}+\frac{\sqrt{3}}{2}i=e^{\frac{\pi}{3}i}$ can not occur. The same argument yields that $A=\frac{1}{2}-\frac{\sqrt{3}}{2}i=e^{-\frac{\pi}{3}i}$ cannot occur. Therefore, $A=-1$, which implies that $\eta=-\frac{1+\sqrt{3}i}{2}$ and
$h(\overline{z})=-h(z)+B=-h(z)+\theta_{1}+\xi_{5},$
@@
Next, we will prove that $L(z)=az+b$ is a linear function with two constant $a(\neq 0),~{}b$. Without loss of generality, we assume that
$h(L(z))=-h(z)+C,$ (2.31)
where $C$ is a constant. Suppose that $h$ is transcendental. We recall the following result (see [11, Theorem 2] and [17, p. 370]).
@@
$\infty=\limsup_{r\not\in E,~{}r\rightarrow\infty}\frac{T(r,h(L(z)))}{T(r,h(z))}=\limsup_{r\not\in E,~{}r\rightarrow\infty}\frac{T(r,-h(z)+c)}{T(r,h(z))}=1,$
@@
$=\frac{1}{m}\frac{1+\epsilon_{1}}{1-\epsilon_{2}}T(r,-h(z)+C))=\frac{1}{m}\frac{1+\epsilon_{1}}{1-\epsilon_{2}}T(r,h(z))+O(1),$
@@
Thus, the above discussion yields that $L(z)=qz+c$ with $q\neq 0$ and $c$ is a constant. So, $h(qz+c)=-h(z)+C$, where $C$ is a constant.
@@
$qh^{\prime}(qz+c)=-h^{\prime}(z).$ (2.33)
If $h^{\prime}(z)$ is a constant, then, $q=-1$ and $|q|=1$. Next, we assume that $h^{\prime}(z)$ is nonconstant. Suppose that $|q|<1$. For each $t^{*}\in\mathbb{C}$, by (2.33), we get
$h^{\prime}(t^{*})=-qh^{\prime}(qt^{*}+c)=(-q)^{2}h^{\prime}[q(qt^{*}+c)+c]=\cdots=(-q)^{n}h^{\prime}[q^{n}t^{*}+c\sum_{k=0}^{n-1}q^{k}].$
@@
$h^{\prime}(t^{*})=h^{\prime}(c\frac{1}{1-q})\lim_{n\rightarrow\infty}(-q)^{n}=0,$
which means that $h^{\prime}(z)\equiv 0$, a contradiction. If $|q|>1$, then we rewrite (2.33) as $h^{\prime}(z)=-ph^{\prime}(pz+c^{\prime})$ with $p=\frac{1}{q}$ and $c^{\prime}=-\frac{c}{q}$. Note that $|p|=|\frac{1}{q}|<1$. The same argument as above yields a contradiction. So, we obtain $|q|=1$.
@@
we assume that one of $\frac{f(z)}{e^{\frac{g(z)}{2}}}$ and $\frac{f(L(z))}{e^{\frac{g(z)}{4}}}$ is not a constant. Via (iii) of Theorem A, we have that $\frac{f(z)}{e^{\frac{g(z)}{2}}}=sn^{\prime}(h(z))$ and $\frac{f(L(z))}{e^{\frac{g(z)}{4}}}=sn(h(z))$, where $h(z)$ is any nonconstnat entire function and $sn$ is Jacobi elliptic function satisfying ${sn^{\prime}}^{2}=1-sn^{4}$. So,

### v2

@@ 1 Introduction and main results
which implies that the above equation admits meromorphic solution $f(z)$. Obviously, $f(z)$ is not the form $de^{\frac{\alpha z+\beta}{3}}$ and the hyper-order of $f$ is $\rho_{2}(f)=1$.
@@
Obviously, all the points in $\mathbb{L}$ are the poles and periods of Weierstrass-$\wp$. Suppose that $D$ is the parallelogram with vertices at 0, $\omega_{1}$, $\omega_{2}$, $\omega_{1}+\omega_{2}$. Note that the order of Weierstrass-$\wp$ is 2. Here, the order is the number of poles of Weierstrass-$\wp$ or the number of zeros of $\wp-a$ ($a\in\mathbb{C}$) in the parallelogram. Together with $(\wp^{\prime})^{2}=4\wp^{3}-1$, one gets that Weierstrass-$\wp$ has two distinct zeros in $D$, say $\theta_{1}$ and $\theta_{2}$. In view of $\wp(z)=0\Leftrightarrow\wp^{\prime}(z)=\pm i$, without loss of generality, throughout the paper, we assume that $\wp^{\prime}(\theta_{1})=-i$ and $\wp^{\prime}(\theta_{2})=i$.
@@
(1). $h(L(z))=Ah(z)+\theta_{1}+\tau_{1}$ with $\tau_{1}\in\mathbb{L}$;
(2). $h(L(z))=Ah(z)+\theta_{2}+\tau_{2}$ with $\tau_{2}\in\mathbb{L}$;
(3). $h(L(z))=Ah(z)+\tau_{3}$ with $\tau_{3}\in\mathbb{L}$, where $A$ is a constant with $A^{3}=-1$ and $A\tau\in\mathbb{L}$ for any period of $\wp(z)$.
@@
(1). $h(z+c)=Ah(z)+\theta_{1}+\tau_{1}$ with $\tau_{1}\in\mathbb{L}$;
(2). $h(z+c)=Ah(z)+\theta_{2}+\tau_{2}$ with $\tau_{2}\in\mathbb{L}$;
(3). $h(z+c)=Ah(z)+\tau_{3}$ with $\tau_{3}\in\mathbb{L}$, where $A$ is a constant with $A^{3}=-1$ and $A\tau\in\mathbb{L}$ for any period of $\wp(z)$.
@@
Example 2. In [39, Theorem 3], Ozawa derived that there exists periodic entire function of arbitrary order $\sigma(\geq 1)$. So, there exists an entire function $g(z)$ such that $\rho(g(z))=\sigma$ and $g(z+c)=g(z)$. Set $h(z)=e^{az}g(z)+b$, where $a$, $b$ are constants with $e^{ac}=A$ and $(1-A)b=\theta_{1}+\tau_{1}$ or $\theta_{2}+\tau_{2}$ or $\tau_{3}$ with $A^{3}=-1$. Then, $\rho(h(z))=\rho(g(z))=\sigma$ and a calculation yields that
$h(z+c)=Ah(z)+(1-A)b.$
@@
if and only if $f(z)=Ae^{\frac{g(z)}{n}}$ and $g(z+c)=\frac{n}{m}g(z)+nLn\frac{B}{A}$, where $A,~{}B$ are two nonzero constants with $A^{n}+B^{m}=1$. In particularly, if the order of $g(z)$ is less than 1, then (1.9) admits nonconstant meromorphic solutions if and only if $n=m$ and $g(z)=\alpha z+\beta$, $f(z)=Ae^{\frac{\alpha z+\beta}{n}}$ with constant $\alpha(\neq 0)$ such that $A^{n}(1+e^{\alpha c})=1$.
@@
Remark 3. From Theorem 3, we see that the equation (1.9) do not admit meromorphic solutions with poles. It is pointed out that the same argument in Theorems 3 can deal with the Fermat-type difference equations $f^{m}(z)+f^{n}(z+c)=e^{g(z)}$ with $n(\geq 2),~{}m(\geq 3)$ and $(m,n)\neq(3,3)$, we omit the details here.
The following corollary follows immediately from Theorem 3.
Corollary 1. Suppose that $P(z)$ is a polynomial and $f(z)$ is a nonconstant meromorphic function. Suppose that $n(\geq 2),~{}m(\geq 2)$ are two integers with $m+n\geq 5$ and $(n,m)\neq(3,3)$. Then, $f(z)$ is a solution of
@@
From Theorem 3 and Corollary 1, one can easily get the following result.
Corollary 2. Suppose that $f(z)$ is a nonconstant meromorphic function, and $n(\geq 2),~{}m(\geq 2)$ are two integers with $m+n\geq 5$. Then, $f(z)$ is a solution of
@@ 2 Proofs of main results
Firstly, we will prove the necessity. Assume that $h(z)$ satisfies one of (1)-(3), we will prove that $f(z)$ is a meromorphic solution of (1.4). Suppose that $h(z)$ satisfies (1). Then $h(L(z))=Ah(z)+\theta_{1}+\tau_{1}$ with $\tau_{1}\in\mathbb{L}$. The assumption that $A\tau\in\mathbb{L}$ for any period of $\wp(z)$ defines a mapping as
$A:\mathbb{L}\rightarrow\mathbb{L},~{}~{}A:\tau\rightarrow A\tau.$
We claim that the mapping $A:$ is a bijection.
Observe that $A=-1$, or $e^{\frac{\pi}{3}i}$, or $e^{-\frac{\pi}{3}i}$. If $A=-1$, clearly, the claim is valid. Suppose that $A=e^{\frac{\pi}{3}i}$. Among all the periods of $\wp(\omega)$, we list the periods which may take the smallest modulus as follows: $\omega_{1}$, $\omega_{1}+\omega_{2}$, $\omega_{2}$, $\omega_{2}-\omega_{1}$, $-\omega_{1}$, $-\omega_{2}-\omega_{1}$, $-\omega_{2}$ and $\omega_{1}-\omega_{2}$. Then, in view of that $A\tau=\tau e^{\frac{\pi}{3}i}$ is also a period of Weierstrass-$\wp$, below, we have two possible cases (see Figures 1-2 below).
Figure 1: (i)
Figure 2: (ii)
(i). $A\omega_{1}=\omega_{2}$ and $A\omega_{2}=\omega_{2}-\omega_{1}$ and $A(\omega_{2}-\omega_{1})=-\omega_{1}$.
(ii). $A\omega_{1}=\omega_{1}+\omega_{2}$ and $A(\omega_{1}+\omega_{2})=\omega_{2}$ and $A\omega_{2}=-\omega_{1}$.
Either (i) or (ii) holds, a simple analysis yields the claim is right. The same argument implies that the claim holds for $A=e^{-\frac{\pi}{3}i}$. Therefore, the claim is proved.
Observe that $A^{3}=-1$, the claim and the definition of $\wp(z)$ as follows
$\wp(z)=\wp\left(z;\omega_{1},\omega_{2}\right):=\frac{1}{z^{2}}+\sum_{\mu,\nu\in\mathbf{Z};\mu^{2}+\nu^{2}\neq 0}\left\{\frac{1}{\left(z+\mu\omega_{1}+\nu\omega_{2}\right)^{2}}-\frac{1}{\left(\mu\omega_{1}+\nu\omega_{2}\right)^{2}}\right\}.$
A calculation yields that
$\wp(Az)$ $=\frac{1}{(Az)^{2}}+\sum_{\mu,\nu\in\mathbf{Z};\mu^{2}+\nu^{2}\neq 0}\left\{\frac{1}{\left(Az+\mu\omega_{1}+\nu\omega_{2}\right)^{2}}-\frac{1}{\left(\mu\omega_{1}+\nu\omega_{2}\right)^{2}}\right\}$ (2.1)
$=\frac{1}{A^{2}}\left\{\frac{1}{z^{2}}+\sum_{\mu,\nu\in\mathbf{Z};\mu^{2}+\nu^{2}\neq 0}\left\{\frac{1}{\{z+\frac{1}{A}(\mu\omega_{1}+\nu\omega_{2})\}^{2}}-\frac{1}{\{\frac{1}{A}(\mu\omega_{1}+\nu\omega_{2})\}^{2}}\right\}\right\}$
$=-A\left\{\frac{1}{z^{2}}+\sum_{\mu,\nu\in\mathbf{Z};\mu^{2}+\nu^{2}\neq 0}\left\{\frac{1}{\{z-A^{2}(\mu\omega_{1}+\nu\omega_{2})\}^{2}}-\frac{1}{\{-A^{2}(\mu\omega_{1}+\nu\omega_{2})\}^{2}}\right\}\right\}$
$=-A\left\{\frac{1}{z^{2}}+\sum_{\mu,\nu\in\mathbf{Z};\mu^{2}+\nu^{2}\neq 0}\left\{\frac{1}{\left(z+\mu\omega_{1}+\nu\omega_{2}\right)^{2}}-\frac{1}{\left(\mu\omega_{1}+\nu\omega_{2}\right)^{2}}\right\}\right\}$
$=-A\wp(z),$
which implies that $\wp^{\prime}(Az)=-\wp^{\prime}(z)$. It is known that
@@
$\wp(Aw+\theta_{1})$ $=\frac{1}{4}[\frac{\wp^{\prime}(Aw)-\wp^{\prime}(\theta_{1})}{\wp(Aw)-\wp(\theta_{1})}]^{2}-\wp(Aw)-\wp(\theta_{1})=\frac{1}{4}[\frac{\wp^{\prime}(Aw)+i}{\wp(Aw)}]^{2}-\wp(Aw)$ (2.2)
$=\frac{1}{4}\frac{(\wp^{\prime}(Aw)+i)^{2}\wp(Aw)}{\wp(Aw)^{3}}-\wp(Aw)=\frac{(\wp^{\prime}(Aw)+i)^{2}\wp(Aw)}{\wp^{\prime}(Aw)^{2}+1}-\wp(Aw)$
$=\frac{(\wp^{\prime}(Aw)+i)^{2}\wp(Aw)}{(\wp^{\prime}(Aw)+i)(\wp^{\prime}(Aw)-i)}-\wp(Aw)=\frac{2i\wp(Aw)}{\wp^{\prime}(Aw)-i}$
$=\frac{-A2i\wp(w)}{-\wp^{\prime}(w)-i}=\frac{A2i\wp(w)}{\wp^{\prime}(w)+i}.$
$\wp^{\prime}(Aw+\theta_{1})$ $=(\frac{2i\wp(w)}{\wp^{\prime}(w)+i})^{\prime}=2i\frac{\wp^{\prime}(w)(\wp^{\prime}(w)+i)-\wp(w)\wp^{\prime\prime}(w)}{(\wp^{\prime}(w)+i)^{2}}$ (2.3)
$=2i\frac{\wp^{\prime}(w)^{2}+i\wp^{\prime}(w)-\wp(w)\wp^{\prime\prime}(w)}{(\wp^{\prime}(w)+i)^{2}}=2i\frac{\wp^{\prime}(w)^{2}+i\wp^{\prime}(w)-\frac{3}{2}(\wp^{\prime}(w)^{2}+1)}{(\wp^{\prime}(w)+i)^{2}}$
$=-i\frac{\wp^{\prime}(w)-3i}{\wp^{\prime}(w)+i}.$
Set $h(z)=\omega$. Then $h(L(z))=Ah(z)+\theta_{1}+\tau_{1}=Aw+\theta_{1}+\tau_{1}$. Note that both Weierstrass-$\wp$ and $\wp^{\prime}$ are elliptic functions with periods $\omega_{1}$ and $\omega_{2}$, we have
$\wp(h(L(z)))=\wp(Aw+\theta_{1}+\tau_{1})=\wp(Aw+\theta_{1}),$
$\wp^{\prime}(h(L(z)))=\wp^{\prime}(Aw+\theta_{1}+\tau_{1})=\wp^{\prime}(Aw+\theta_{1}).$
@@
$f(L(z))$ $=\frac{1}{2}\frac{1+\frac{\wp^{\prime}(h(L(z)))}{\sqrt{3}}}{\wp(h(L(z)))}=\frac{1}{2}\frac{1+\frac{\wp^{\prime}(Aw+\theta_{1})}{\sqrt{3}}}{\wp(Aw+\theta_{1})}=\frac{1}{2}\frac{1-\frac{1}{\sqrt{3}}i\frac{\wp^{\prime}(w)-3i}{\wp^{\prime}(w)+i}}{\frac{A2i\wp(w)}{\wp^{\prime}(w)+i}}$ (2.4)
$=\frac{1}{2}\frac{(\wp^{\prime}(w)+i)-\frac{1}{\sqrt{3}}i(\wp^{\prime}(w)-3i)}{A2i\wp(w)}=\frac{1}{2}\frac{-1-\sqrt{3}i}{2}\frac{1}{-A}\frac{1-\frac{\wp^{\prime}(w)}{\sqrt{3}}}{\wp(w)}=\frac{1}{-A}\frac{\eta}{2}\frac{1-\frac{\wp^{\prime}(w)}{\sqrt{3}}}{\wp(w)},$
where $\eta=\frac{-1-\sqrt{3}i}{2}$ and $\eta^{3}=1$. Thus, by (i) of Theorem A, one can easily derive that $f(z)^{3}+f(L(z))^{3}=1$ and $f(z)$ is a solution of (1.4). If $h(z)$ satisfies (2) or (3), then the same argument yields that $f(z)$ is also a solution of (1.4).
@@
Observe that $A=\frac{(1+i\frac{\sqrt{3}}{3})}{\eta(1-i\frac{\sqrt{3}}{3})}$ and $\eta^{3}=1$. A calculation yields $A^{3}=-1$ and
@@
Meanwhile, (2.29) yields that
$h(\overline{z})=Ah(z)+B=Ah(z)+\theta_{1}+\xi_{5},$
@@
Next, we will prove that $L(z)$ is a linear function with two constant $a(\neq 0),~{}b$. Without loss of generality, we assume that
$h(L(z))=Ah(z)+C,$ (2.32)
where $A^{3}=-1$ and $C$ is a constant. Suppose that $h$ is transcendental. We recall the following result (see [11, Theorem 2] and [17, p. 370]).
@@
$\infty=\limsup_{r\not\in E,~{}r\rightarrow\infty}\frac{T(r,h(L(z)))}{T(r,h(z))}=\limsup_{r\not\in E,~{}r\rightarrow\infty}\frac{T(r,Ah(z)+C)}{T(r,h(z))}=1,$
@@
$=\frac{1}{m}\frac{1+\epsilon_{1}}{1-\epsilon_{2}}T(r,Ah(z)+C))=\frac{1}{m}\frac{1+\epsilon_{1}}{1-\epsilon_{2}}T(r,h(z))+O(1),$
@@
Thus, the above discussion yields that $L(z)=qz+c$ with $q\neq 0$ and $c$ is a constant. So, $h(qz+c)=Ah(z)+C$, where $C$ is a constant.
@@
$qh^{\prime}(qz+c)=Ah^{\prime}(z).$ (2.34)
If $h^{\prime}(z)$ is a constant, then, $q=A$ and $|q|=|A|=1$. Next, we assume that $h^{\prime}(z)$ is nonconstant. Suppose that $|q|<1$. For each $t^{*}\in\mathbb{C}$, by (2.34), we get
$h^{\prime}(t^{*})=\frac{q}{A}h^{\prime}(qt^{*}+c)=(\frac{q}{A})^{2}h^{\prime}[q(qt^{*}+c)+c]=\cdots=(\frac{q}{A})^{n}h^{\prime}[q^{n}t^{*}+c\sum_{k=0}^{n-1}q^{k}].$
@@
$h^{\prime}(t^{*})=h^{\prime}(c\frac{1}{1-q})\lim_{n\rightarrow\infty}(\frac{q}{A})^{n}=0,$
which means that $h^{\prime}(z)\equiv 0$, a contradiction. If $|q|>1$, then we rewrite (2.34) as $h^{\prime}(z)=Aph^{\prime}(pz+c^{\prime})$ with $p=\frac{1}{q}$ and $c^{\prime}=-\frac{c}{q}$. Note that $|p|=|\frac{1}{q}|<1$. The same argument as above yields a contradiction. So, we obtain $|q|=1$.
@@
we assume that one of $\frac{f(z)}{e^{\frac{g(z)}{2}}}$ and $\frac{f(L(z))}{e^{\frac{g(z)}{4}}}$ is not a constant. Via (iii) of Theorem A, we have that $\frac{f(z)}{e^{\frac{g(z)}{2}}}=sn^{\prime}(h(z))$ and $\frac{f(L(z))}{e^{\frac{g(z)}{4}}}=sn(h(z))$, where $h(z)$ is any nonconstant entire function and $sn$ is Jacobi elliptic function satisfying ${sn^{\prime}}^{2}=1-sn^{4}$. So,